Find the expected value, μ, and standard deviation, σ, for a binomial random variable with each of the following values of n and p. (Round all answers for σ to four decimal places.) (a) n = 50, p = 1/2. μ = σ =

(b) n = 300, p = 1/4. μ = σ =

(c) n = 1000, p = 1/5. μ = σ =

(d) n = 1, p = 0.3. μ = σ =

(e) n = 20, p = 2/5. μ = σ =
Answer #1

Ans:

We know that

mean,\mu=np

standard deviation,\sigma=sqrt(np(1-p))

e.g. for a)n=50,p=0.5

mean=np=50*0.5=25

standard deviation=sqrt(50*0.5*0.5)=3.5355

n p mean std. dev.
a 50 0.5 25 3.5355
b 300 0.25 75 7.5000
c 1000 0.2 200 12.6491
d 1 0.3 0.3 0.4583
e 20 0.4 8 2.1909